OpenStudy (anonymous):

0,7,26,63,124, fimd the missimg number and the explanation

OpenStudy (anonymous):

One pattern that I see is that the numbers are all primes

OpenStudy (zzr0ck3r):

hmm the difference is always prime 7 19 37 61 so maybe that

OpenStudy (zzr0ck3r):

ha

OpenStudy (anonymous):

yeah, what @zzr0ck3r said

OpenStudy (anonymous):

series x^3-1

OpenStudy (anonymous):

huh?

OpenStudy (anonymous):

1^3-1,2^2-1,3^3-1

OpenStudy (zzr0ck3r):

yeah

OpenStudy (anonymous):

I think it is the addition of prime numbers that gives the next number in the sequence

OpenStudy (zzr0ck3r):

for some reason I read 19 was in there....

OpenStudy (anonymous):

That would be the most logical answer

OpenStudy (zzr0ck3r):

nah its \(n^3-1\)

OpenStudy (anonymous):

\[x ^{3}-1 where x=1,2,3,.....\]

OpenStudy (anonymous):

7, 19, 37... etc

OpenStudy (anonymous):

but there is a problem with that because the the next prime number after 7 is 11 so you are right, that doesn't work

OpenStudy (zzr0ck3r):

\(\{n^3-1 \ | \ n\in \mathbb{N}\}\)

OpenStudy (anonymous):

I'm not even sure what that is

OpenStudy (anonymous):

what is n^3 - 1? is that the cube of a number and subtract 1?

OpenStudy (zzr0ck3r):

lol the answer is 215

OpenStudy (zzr0ck3r):

yes

OpenStudy (zzr0ck3r):

1^3-1 = 0 2^3-1 = 7 3^3-1 = 26 4^3-1 = 63 5^3-1 = 124 6^3-1 = 215

OpenStudy (anonymous):

oh I got it, good problem